# SimCity: Multi-Agent Urban Development Simulation with Rich Interactions

### v2

@@ 6 Limitations
Our current simulation framework does not incorporate real-world financial markets, such as those for bonds, stocks, or their derivatives. Moreover, the simulated agents do not employ mathematical optimization techniques in their investing strategies. These simplifications may limit the model’s ability to accurately simulate complex market dynamics.
@@ 7 Conclusion
With SimCity, we conduct complex simulations that incorporate four distinct LLM-driven economic roles (households, firms, the central bank, and the government), with a visualized map and markets with heterogeneous goods. Experiments show that SimCity robustly matches a checklist of established macroeconomic phenomena and exhibits plausible responses to exogenous shocks. We hope SimCity could serve as a solid base platform for a more realistic economic simulation.
Acknowledgement
We sincerely thank Jian Li and Kaifeng Lyu for their valuable discussions and insightful comments that help improve this work.
@@ 8 Ethics Statement
9 Reproducibility Statement
To help with the reproducibility of our work, we provide all LLMs prompts in Appendix E.1. We call LLMs API with default sampling parameters. We repeat experiments with different random seeds (Appendix B.1) to ensure that our main results are robust. We plan to release code in the near future.
@@ B.1 Robustness
We run three sets of experiments with the same hyper-parameters but different random seeds, and the results demonstrate that the observed regularity is reproducible.
@@ B.2 Price Impulse
Figure 8: External impulse does not significantly affect long-run prices.
@@ B.3 Ablation Study: Visual Module
As illustrated in Figure 9, the absence of the visual module significantly degrades the spatial logic of urban planning. While the VLM-based approach (Figure 9a) fosters a natural, concentric city structure with a distinct center, the text-only approach (Figure 9b) fails to interpret spatial density effectively. Consequently, the agents form an artificial linear arrangement of buildings (”two lines”), confirming that the visual module is crucial for interpreting complex spatial information that is difficult to convey through text alone.
@@ C.2 Households
  • Heterogeneous Consumption Preferences: The household has needs denoted by $n_{i}$. For each good $g\in\mathcal{G}$, $n_{i,g}$ represents the desired number of units of good $g$ to be purchased. $n_{i}$ may split to essential needs $n_{e,i}$ which is fixed denoting immutable demands and additional needs $n_{a,i}$. The modeling of essential goods has a long history in economics. See for example, the Stone-Geary Utility function (Stone, 1954).
@@ D.1 Environment Setup
Initial prices of all goods are set at 50. However, the specific choice of initial prices do not matter for our simulation. We experimented with various initial prices but in all cases, as we mention in 5.4, prices reach steady-state in a small number of steps. Similarly, wages of positions and rents of residential buildings are set arbitrarily.
@@ D.4 Profile Setup
Age distribution is from the Demographic and Housing Characteristics (DHC) table from U.S. Census Bureau [^8] https://data.census.gov/table?q=PCT12&d=DEC+Demographic+and+Housing+Characteristics.
@@ D.5 Synthesis
We assume that each category of goods is produced by a representative type of firm. Therefore, we generate a firm template for each category, which specifies the firm’s name, the skill requirements for its positions, and an input-output ”recipe” derived from the Input-Output Tables (IOTable) provided by the OECD. [^10] https://www.oecd.org/en/data/datasets/input-output-tables.html
@@
Observing that the positions required skills in the generated firm templates contain many semantically similar skills, we use the prompt introduced in Listing 2 to call an LLM to merge these similar skills. Skills after merge are shown at Table 3.
@@ E.1 Prompting Structure
As mentioned in 4, the Agent will receive the necessary information to make its decisions. The Profile includes some basic personal information about the Agent, such as current income, age, skills, job, and residence. The Report includes some of the Agent’s recent experiences and information. The Observation presents information about the overall environment that the Agent should be aware of, such as vacant positions, the unemployment rate, prices of goods and the Return on Investment (ROI).
@@ E.2 Function Calling
The model interacts with the environment by means of Function Calling. The framework loads all the operations that the Agent can execute and appends formatted function names along with their descriptions to the prompt. The LLMs will return the actions to be taken and their parameters in JSON format. Framework will execute after a verification.

### v3

@@ 6 Limitations
Our current simulation framework does not incorporate real-world financial markets, such as those for bonds, stocks, or their derivatives. Moreover, the simulated agents do not employ mathematical optimization techniques in their investing strategies. These simplifications may limit the model’s ability to simulate complex market dynamics accurately.
@@ 7 Conclusion
With SimCity, we conduct complex simulations that incorporate four distinct LLM-driven economic roles (households, firms, the central bank, and the government), with a visualized map and markets with heterogeneous goods. Experiments show that SimCity robustly matches a checklist of established macroeconomic phenomena. Moreover, when applied to novel economic scenarios, SimCity generates economically coherent and informative predictions. We hope SimCity could serve as a solid base platform for a more realistic economic simulation.
@@ 8 Ethics Statement
Acknowledgement
We sincerely thank Jian Li, Kaifeng Lyu, and Boyuan Zheng for their valuable discussions and insightful comments that help improve this work.
@@ B.1 Robustness
We run three sets of experiments with the same hyperparameters but different random seeds, and the results demonstrate that the observed regularity is reproducible.
@@
Furthermore, we evaluate SimCity using gpt-5 as the reasoning backbone. Figure 9 shows that the fundamental relationships remain consistent under a more advanced LLM.
Figure 9: Macroeconomic patterns simulated using the gpt-5 driven agents.
@@ B.2 Price Impulse
Figure 10: External impulse does not significantly affect long-run price trends.
@@ B.3 Ablation Study: Visual Module
As illustrated in Figure 11, the absence of the visual module significantly degrades the spatial logic of urban planning. While the VLM-based approach (Figure 11a) fosters a natural, concentric city structure with a distinct center, the text-only approach (Figure 11b) fails to interpret spatial density effectively. Consequently, the agents form an artificial linear arrangement of buildings ("two lines"), confirming that the visual module is crucial for interpreting complex spatial information that is difficult to convey through text alone.
@@
B.4 Extra Example for City Expansion
An extra example for city expansion, Figure 12, is shown here.
Figure 12: Another example of city dynamics during the move-in phase.
@@ C.2 Households
  • Heterogeneous Consumption Preferences: The household has needs denoted by $n_{i}$. For each good $g\in\mathcal{G}$, $n_{i,g}$ represents the desired number of units of good $g$ to be purchased. $n_{i}$ may split to essential needs $n_{e,i}$ which is fixed denoting immutable demands and additional needs $n_{a,i}$. The modeling of essential goods has a long history in economics. See, for example, the Stone-Geary Utility function (Stone, 1954).
@@ D.1 Environment Setup
Initial prices of all goods are set at 50. However, the specific choice of initial prices does not matter for our simulation. We experimented with various initial prices, but in all cases, as we mention in 5.5, prices reach steady-state in a small number of steps. Similarly, wages of positions and rents of residential buildings are set arbitrarily.
@@ D.4 Profile Setup
Age distribution is from the Demographic and Housing Characteristics (DHC) table from the U.S. Census Bureau [^8] https://data.census.gov/table?q=PCT12&d=DEC+Demographic+and+Housing+Characteristics.
@@ D.5 Synthesis
We assume that each category of goods is produced by a representative type of firm. Therefore, we generate a firm template for each category, which specifies the firm’s name, the skill requirements for its positions, and an input-output "recipe" derived from the Input-Output Tables (IOTable) provided by the OECD. [^10] https://www.oecd.org/en/data/datasets/input-output-tables.html
@@
Observing that the positions required skills in the generated firm templates contain many semantically similar skills, we use the prompt introduced in Listing 2 to call an LLM to merge these similar skills. Skills after merge are shown in Table 3.
@@ E.1 Prompting Structure
As mentioned in Section 4, the Agent will receive the necessary information to make its decisions. The Profile includes some basic personal information about the Agent, such as current income, age, skills, job, and residence. The Report includes some of the Agent’s recent experiences and information. The Observation presents information about the overall environment that the Agent should be aware of, such as vacant positions, the unemployment rate, prices of goods, and the Return on Investment (ROI).
@@ E.2 Function Calling
The model interacts with the environment by means of Function Calling. The framework loads all the operations that the Agent can execute and appends formatted function names along with their descriptions to the prompt. The LLMs will return the actions to be taken and their parameters in JSON format. The framework will execute after a verification.